OpenStudy (anonymous):

solve x 7/8 -1/x = 3/4

OpenStudy (anonymous):

move the 7/8 over \[\frac{-1}{x}=\frac{-7}{8}+\frac{6}{8}=\frac{-1}{8}\]

OpenStudy (anonymous):

solve x